# Chiral cat state generation via the Sagnac-Fizeau effect

### v1

@@ Abstract
The Schrödinger cat states are a kind of significant quantum resource in quantum physics and quantum information science. Here we propose a proposal for generating chiral cat states in a spinning resonator supporting both the clockwise (CW) and counterclockwise (CCW) traveling modes, which are dispersively coupled to a two-level atom. The physical mechanism for the chiral cat-state generation is based on the Sagnac-Fizeau effect. Concretely, when the resonator is rotating, the CW and CCW modes have different frequency detuning with respect to the atomic transition frequency, and hence the atomic-state-dependent rotating angular velocities for the CW and CCW modes in phase space are different. This mode-dependent evolution leads to the chirality in the state generation. By choosing proper system parameters, we achieve the separate generation of the Schrödinger cat states and coherent states in the two traveling modes. We also investigate quantum coherence properties of the generated states by examining their Wigner functions. In addition, the influence of the system dissipations on the state generation in the open-system case is investigated. Our work will provide some insights into the development of chiral optical devices and nonreciprocal photonics physics.
@@ I INTRODUCTION
The Schrödinger cat states are an essential resource for studying fundamental quantum physics Schroedinger1935 ; Peres2002 , including exploring macroscopic quantum coherence LeggettPRL1985 ; LeggettJCM2002 , quantum decoherence ZurekRMP2003 ; Joos2003 ; BrunePRL1996 , and quantum-classical boundary ZurekPT1991 ; HarocheRMP2013 . They also have wide application potential in quantum information processing, such as quantum teleportation LiaoPLA2006 ; FurusawaSci1998 ; BraunsteinPRL1998 , quantum communication JeongPRA2001 ; BraunsteinRMP2005 ; VlastakisSc2013 ; Gilchrist2004 , quantum computation CochranePRA1999 ; RalphPRA2003 ; MirrahimiNJP2014 ; Ofekature2016 ; Guillaud2023 ; JeongPRA2002 , and quantum precision measurement JooPRL2011 ; MunroPRA2002 ; BildSci2023 . So far, various methods have been proposed to create the Schrödinger cat states, including conditional rotation mechanism VlastakisSc2013 ; GuoPLA1996 ; BrunePRA1992 , conditional displacement mechanism MonroeSciece1996 ; BosePRA1997 ; ArmourPRL2002 ; AshhabPRA2010 ; LiaoPRL2016 ; LiaoPRA2016 ; ChenPRL2021 , non-Gaussian measurement OurjoumtsevNa2007 ; SunPRL2021 ; HuPRR2023 ; OurjoumtsevScien2006 ; TakahashiPRL2008 ; HanLPR2023 , quantum state transfer HoffPRL2016 ; TehPRA2018 , Kerr-nonlinear evolution MiranowiczQO1990 ; KuangPRA2003 ; LiaoPRA2020 ; GrimmNa2020 ; YinPRA2022 ; HeNC2023 , and reservoir engineering techniques PoyatosPRL1996 ; AsjadASjas2014 ; BrunelliPRA2018 . However, most previous schemes have seldomly focused on designing the cat states with directional differences LodahlA2017 , making them less suitable for nonreciprocal applications such as nonreciprocal quantum communications FengSC2011 ; ScheucherSC2016 , chiral optical devices LodahlA2017 ; SounasNP2017 ; TangPRL2022 ; ZuoOE2024 , and stable codes for quantum error correction TerhalRMP2015 ; GeorgescuNEP2020 . As a result, how to create chiral cat states remains an interesting and desired topic in quantum physics and quantum optics.
@@
Quantum nonreciprocal devices are fundamental components in chiral quantum technologies, and much recent attention has been paid to nonreciprocal physics SounasNP2017 ; YangPRL2019 ; JiaoPRL2020 ; DongSC2021 ; XuPRAP2020 ; XuPRA2021 ; MetelmannPRX2015 ; FruchartNa2021 . Spinning resonators, which exhibit the broken time-reversal symmetry, provide an ideal platform for realizing nonreciprocal devices MaayaniNa2018 ; MalykinPU2000 ; JingOpti2018 . Recently, some significant progress has been achieved based on spinning resonators, including nonreciprocal photon blockade HuangPRL2018 , nonreciprocal entanglement JiaoPRL2020 ; JiaoPRAA2022 , and nonreciprocal topological phonon transfer LaiPRL2024 . However, the in-depth exploration of chirality-related mechanisms in spinning resonators, particularly the controlled generation and manipulation of chiral quantum states, remains an ongoing challenge. Addressing these issues is crucial for both understanding and practical applications of chiral quantum technologies. Note that the nonreciprocal generation of the Schrödinger cat states induced by the topological effect has been theoretically studied. LiSC2024 .
@@
Inspired by both the Schrödinger-cat-state generation via the dispersive Jaynes-Cummings model GuoPLA1996 and the Sagnac-Fizeau effect MalykinPU2000 , we propose an approach for creating chiral cat states in a system comprising a two-level atom coupled to a spinning resonator, which supports both clockwise (CW) and counterclockwise (CCW) traveling modes. The rotation of the resonator induces a Sagnac-Fizeau effect, resulting in different effective frequencies for the CW and CCW modes relative to the atomic frequency. The effective differences in frequency detunings of the CW and CCW modes induce distinct quantum state generation for the two traveling modes. Concretely, by adjusting the system parameters, we prepare a coherent state in one of the traveling modes and a cat state in the other one. We investigate the coherence properties of these generated states by examining their Wigner functions. We also consider the influence of the system dissipations on the state generation. Our proposal paves the way for developing quantum chiral devices based on the coupled atom-cavity systems.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the physical model and derive the effective Hamiltonians. In Sec. III, we investigate the generation of chiral cat states and analyze the nonclassical properties of these generated states by examining their Wigner functions. We also numerically calculate the exact state generation and verify the validity of the approximation with the fidelity. In Sec. IV, we present numerical simulations of the state generation in the open-system case, and investigate the influence of the system dissipations on the state generation. Finally, we discuss the experimental implementation of this scheme and conclude this work in Sec. V.
@@ II Physical model and Hamiltonians
We consider a cavity-QED system consisting of a two-level atom with an energy separation $\omega_{a}$ coupled to a spinning resonator supporting both the CW and CCW traveling modes via the optical evanescent fields (Fig. 1). The rotation of the resonator in the CCW direction at an angular velocity $\Omega$ induces a Sagnac-Fizeau frequency shift for both the CW and CCW modes. Then the effective frequencies of the CW and CCW modes become $\omega_{c}+\Delta_{\mathrm{sag}}$ and $\omega_{c}-\Delta_{\mathrm{sag}}$, respectively. Here, $\omega_{c}$ is the resonance frequency of the stationary resonator, and $\Delta_{\mathrm{sag}}$ is the Sagnac-Fizeau frequency shift defined as MalykinPU2000
@@
In Eq. (1), $\lambda$ and $c$ represent the wavelength of the field mode and the speed of light in vacuum, respectively, $n_{r}$ is the refractive index, and $R$ is the radius of the resonator. The term $dn_{r}/d\lambda$ is the dispersion term related to the relativistic origin of the Sagnac effect, which is relatively small and can be ignored in typical materials MaayaniNa2018 ; MalykinPU2000 .
@@
where $\sigma_{z}\equiv|e\rangle\langle e|-|g\rangle\langle g|$ is $z$-direction Pauli operator, with $|g\rangle$ and $|e\rangle$ being the ground and excited states of the two-level atom, respectively. The operators $\sigma_{-}\equiv|g\rangle\langle e|$ and $\sigma_{+}\equiv|e\rangle\langle g|$ are, respectively, the lowering and raising operators of the atom. The operators $a_{\epsilon}^{\dagger}$ and $a_{\epsilon}$ ($\epsilon=\mathrm{CW},\mathrm{CCW}$) are the creation and annihilation operators of the $\epsilon$ mode. The variables $\Delta_{\mathrm{CW}}\equiv\Delta+\Delta_{\mathrm{sag}}$ and $\Delta_{\mathrm{CCW}}\equiv\Delta-\Delta_{\mathrm{sag}}$ represent, respectively, the effective frequency detunings of the CW and CCW modes with respect to the resonance frequency of the atom, with $\Delta\equiv\omega_{c}-\omega_{a}$. The parameter $J$ denotes the coupling strength between the two traveling modes and the atom.
In this work, we consider the dispersive regime of the cavity-field-atom couplings, namely, the detunings between the two-level atom and the cavity field are much larger than the coupling strengths, i.e., $\Delta_{\epsilon=\mathrm{CW},\mathrm{CCW}}\gg J\sqrt{n_{\epsilon}+1}$ Schleich2001 , where $n_{\epsilon}$ are the maximally involved photon numbers in the two cavity modes. In this case, we can perform the Frohlich-Nakajima transformation FPR1950 ; Nakajima1955 on the Hamiltonian (2) to derive the effective Hamiltonian by eliminating the coherent transitions between the two atomic states,
@@
$\zeta_{\epsilon=\mathrm{CW},\mathrm{CCW}}/4\Delta_{\mathrm{sag}}\ll 1,$ (4)
@@
The physical mechanism of the chiral cat-state generation can be understood more clearly by expressing the approximate Hamiltonian of the system in the form of conditional rotation couplings
@@
Here, the opposite signs of the coefficients for $|e\rangle\langle e|$ and $|g\rangle\langle g|$ can be used to generate cat states for the traveling modes, and the different values of $\zeta_{\mathrm{CW}}$ and $\zeta_{\mathrm{CCW}}$ lead to different phase space positions for the two traveling modes, since the effective rotating velocities of the two modes are different.
III Generation of the chiral
cat states
@@ III Generation of the chiral
with the normalization coefficients
@@
Typically, the states $|\psi_{\pm}(t)\rangle$ of the CW and CCW modes are entangled states. However, when the two coherent state amplitudes for one of the two modes are overlapped, i.e., $\alpha^{(+)}_{\mathrm{CW}/\mathrm{CCW}}=\alpha^{(-)}_{\mathrm{CW}/\mathrm{CCW}}$, then the two modes will be disentangled. For example, we set the effective detunings $\Delta_{\mathrm{CW}}=2\Delta_{\mathrm{CCW}}$, then the CW mode is in a maximally distinguishable coherent state (cat state) while the CCW mode can remain in a coherent state at time $t_{s}=\pi/2\zeta_{\mathrm{CW}}$. We denote this phenomenon as the generation of chiral cat states, i.e., creating distinct states in the two different rotating modes of the cavity. When $\Delta_{\mathrm{CW}}\neq 2\Delta_{\mathrm{CCW}}$, it becomes impossible to simultaneously achieve a coherent state and a maximally distinguishable cat state. In this case, one mode is in a coherent state while the other mode is in a less distinguishable cat state. Nonetheless, a chiral cat state can still be generated under these conditions.
Figure 2: The Wigner functions $\mathrm{W}_{\mathrm{CW,CCW}}^{(\pm)}(\chi)$ of the generated states $\rho^{(\pm)}_{\mathrm{CW,CCW}}$ at time $t_{s}=\pi/2\zeta_{\mathrm{CW}}$. Panels (a) and (b) [(c) and (d)] show the Wigner functions $\mathrm{W}_{\mathrm{CW}}^{(+)}(\chi)$ and $\mathrm{W}_{\mathrm{CW}}^{(-)}(\chi)$ [$\mathrm{W}_{\mathrm{CCW}}^{(+)}(\chi)$ and $\mathrm{W}_{\mathrm{CCW}}^{(-)}(\chi)$] for the CW (CCW) mode, respectively. The coherent amplitude of the initial coherent state $|\alpha\rangle$ is $\alpha=1.8$. Other parameters used are $\Delta/J=33$ and $\Delta_{\mathrm{sag}}/J=11$.
To demonstrate the quantum coherence effect in the generated chiral cat states, we calculate the Wigner function Barnett1997 for both the CW and CCW modes. When the CW and CCW modes are in the states described by the density matrice $\rho_{\mathrm{CW}}$ and $\rho_{\mathrm{CCW}}$, respectively, the Wigner functions are defined by
@@
In Figs. 2(a)-2(d), we show the Wigner functions $\mathrm{W}_{\mathrm{CW,CCW}}^{(\pm)}(\chi)$ for both the CW and CCW modes at time $t_{s}=\pi/2\zeta_{\mathrm{CW}}=69.11J^{-1}$. Concretely, Figs. 2(a) and 2(b) display the Wigner functions $\mathrm{W}_{\mathrm{CW}}^{(+)}(\chi)$ and $\mathrm{W}_{\mathrm{CW}}^{(-)}(\chi)$ for the CW mode, respectively. Here we can see distinct interference fringes caused by quantum superposition of two coherent states. The interference fringes indicate the presence of coherent superposition states, commonly referred to as cat states. In contrast, the Wigner functions for the CCW mode exhibiting Gaussian peaks in the phase space, as shown in Figs. 2(c) and 2(d), which are signature for coherent states in the CCW mode. We note that the appearance of the coherent state for the CCW mode is caused by the overlap between the two coherent amplitudes $\alpha_{\mathrm{CCW}}^{(-)}(t_{s})$ and $\alpha_{\mathrm{CCW}}^{(+)}(t_{s})$. Comparing Figs. 2(a) [2(b)] with 2(c) [2(d)], we observe that the CW mode is in a cat state while the CCW mode is in a coherent state, demonstrating the generation of chiral cat states.
@@
Figure 3: The analytical measurement probabilities $\mathcal{P}_{\pm}(t)$ (a) from Eq. (12) and the exact measurement probabilities $P_{\pm}(t)$ (b) from Eq. (19) as functions of the scaled evolution time $Jt$. For the exact solution in panel (b), both the CW and CCW modes have a truncation dimension of 13. Other parameters used in panels (a) and (b) are the same as those in Fig. 2.
In Fig. 3(a), we display the analytical measurement probabilities $\mathcal{P}_{\pm}(t)$, given by Eq. (12), vs the scaled evolution time $Jt$. We observe that the sum of $\mathcal{P}_{+}(t)$ and $\mathcal{P}_{-}(t)$ is normalized. Furthermore, for longer subsequent measurement time $t_{m}$, the measurement probabilities $\mathcal{P}_{\pm}(t)\approx 1/2$, corresponding to $\sum_{l=1,-1}\exp[2li\vartheta(t_{m})+\left|\alpha\right|^{2}[\sum_{\epsilon=\mathrm{CW},\mathrm{CCW}}\mathrm{exp}(2li\zeta_{\epsilon}t_{m})-2]]\rightarrow 0$ in Eq. (12). The exact measurement probabilities $P_{\pm}(t)$ given by Eq. (19) are shown in Fig. 3(b). The exact measurement probabilities $P_{\pm}(t)$ exhibit oscillating behavior around 1/2, which is caused by the last term in Eq. (3). By comparing the analytical results with the numerical results, we find that the analytical probability evolution is smooth, while the numerical results exhibit oscillation. Nevertheless, the pattern of the approximate and exact results agree well, which confirms the validity of the approximate made during the derivation of the approximate Hamiltonian.
To quantitatively verify the validity of the approximate Hamiltonian $H_{\mathrm{app}}$, we evaluate the fidelity $F(t)=|\langle\Psi(t)|\psi_{\mathrm{app}}(t)\rangle|^{2}$ between the exact and approximate states. Here, $|\psi_{\mathrm{app}}(t)\rangle$ represents the approximate state governed by the approximate Hamiltonian $H_{\mathrm{app}}$ in Eq. (5), and $|\Psi(t)\rangle$ represents the exact state governed by the full Hamiltonian $H$ in Eq. (2). Based on Eqs. (2) and (5), the fidelity can be calculated as
@@
Figure 4: The fidelities (a) $F(t)$, (b) $F_{+}(t)$, and (c) $F_{-}(t)$ as functions of the scaled evolution time $Jt$. The parameters used are the same as those in Fig. 2.
Similarly, we can also evaluate the performance of the chiral cat-states generation by calculating the fidelities between the generate states $|\Psi_{\pm}(t)\rangle$ in Eq. (18) and the target states $|\psi_{\pm}(t)\rangle$ in Eq. (10). Using Eqs. (18) and (10), the fidelities $F_{\pm}(t)=|\langle\Psi_{\pm}(t)|\psi_{\pm}(t)\rangle|^{2}$ can be obtained as
@@
III.3 The Wigner functions of the exact states
Figure 5: (a)-(d) The Wigner functions $\mathcal{W}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ for the exact states $|\Psi_{\pm}(t_{s})\rangle$ in Eq. (18) at time $t_{s}=\pi/2\zeta_{\mathrm{CW}}$. Other parameters used are the same as those in Fig. 2.

### v2

# Chiral-cat-state generation via the Sagnac-Fizeau effect

@@ Abstract
Chiral quantum state generation is an interesting topic in quantum physics and quantum information science. Here we propose an approach for generating chiral cat states in a spinning resonator supporting both the clockwise (CW) and counterclockwise (CCW) traveling modes, which are dispersively coupled to a two-level atom. The physical mechanism for the chiral-cat-state generation is based on the Sagnac-Fizeau effect. Concretely, when the resonator is rotating, the CW and CCW modes have different frequency detuning with respect to the atomic transition frequency and hence the atomic-state-dependent rotating angular velocities for the CW and CCW modes in phase space are different. This mode-dependent evolution leads to a chirality mechanism in the state generation. Based on the mode-dependent conditional rotation evolution and atomic projection measurement at proper time, we achieve the separate generation of cat states in the CW mode and coherent states in the CCW mode. We also investigate quantum coherence properties of the generated states by examining their Wigner functions. In addition, the influence of the system dissipations on the state generation in the open-system case is investigated. Our work will provide some insights into the development of chiral optical devices and nonreciprocal photonics.
@@ I   INTRODUCTION
The Schrödinger cat states are an essential resource for studying fundamental quantum physics [1, 2], including exploring macroscopic quantum coherence [3, 4], quantum decoherence [5, 6, 7], and quantum-classical boundary [8, 9]. They also have wide application potential in quantum information processing, such as quantum teleportation [12, 11, 10], quantum communication [14, 13, 16, 15], quantum computation [17, 19, 20, 21, 22, 18], and quantum precision measurement [24, 23, 25]. So far, various methods have been proposed to create the Schrödinger cat states, including conditional rotation mechanism [16, 27, 26], conditional displacement mechanism [28, 29, 30, 31, 32, 33, 34], non-Gaussian measurement [36, 38, 39, 35, 37, 40], quantum state transfer [41, 42], Kerr-nonlinear evolution [43, 44, 45, 46, 47, 48], and reservoir engineering techniques [49, 50, 51]. Motivated by the rapid development of chiral and nonreciprocal quantum devices, it becomes an interesting topic to prepare chiral quantum resources [52, 53], including chiral quantum superposition, entanglement, and squeezing. These resources are important elements for implementing nonreciprocal application such as nonreciprocal quantum communications [54, 55], chiral optical devices [56, 57, 58], and stable codes for quantum error correction [59, 60]. Consequently, how to create chiral cat states, which exhibit mirror asymmetry in quantum superposition, remains an interesting and desired task in chiral quantum optics [52, 53] and nonreciprocal photonics [56]. Note that the chirality, which is based on the spatial reflection asymmetry [61], has been suggested to implement nonreciprocal quantum devices [62, 63], in which the nonreciprocity is usually introduced by the breaking of the time-reversal symmetry [61].
@@
Quantum nonreciprocal devices are fundamental components in chiral quantum technologies and much recent attention has been paid to nonreciprocal physics [65, 66, 68, 67, 70, 64, 69]. The spinning resonators [72, 71, 73], as a typical nonreciprocal device [65, 66, 68, 67, 70, 64, 69, 72, 71, 73], could be used to provide an ideal platform for studying chiral quantum phenomena due to the equivalence between the spatial reflection and the time-reversal operation in this system. Recently, significant progress has been achieved based on spinning resonators, including nonreciprocal photon blockade [74], nonreciprocal entanglement [66, 75], and nonreciprocal topological phonon transfer [76]. However, the controlled generation and manipulation of chiral quantum states remain to be explored. Addressing these issues is crucial for both understanding and practical applications of chiral quantum technologies.
@@
In this paper, we explore the generation of chiral cat states in a bichromatic Jaynes-Cummings system [77, 78, 79] comprising a two-level atom coupled to a spinning resonator, which supports both clockwise (CW) and counterclockwise (CCW) traveling modes. The rotation of the resonator induces a Sagnac-Fizeau effect [71], resulting in different effective detunings for the CW and CCW modes relative to the atomic frequency. We consider the large detuning case between the atom and the two cavity modes, which enables a dispersive interaction. The different frequency detunings of the CW and CCW modes will induce different conditional rotation in phase space for the two traveling modes, which can be used to provide a chirality mechanism for generating distinct states. Concretely, by performing a projection measurement on the atom at a specific time, a cat state can be prepared in the CW mode and a coherent state in the CCW mode, or different cat states can be generated in both modes at other measurement times. Here, the simultaneous preparation of different quantum states in the CW and CCW modes of the spinning resonator is referred to as chiral quantum state generation. We further investigate the coherence properties of these generated states by examining their Wigner functions. We also consider the influence of the system dissipations on the state generation. Our proposal paves the way for developing quantum chiral devices based on the coupled atom-cavity systems.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce the physical model and derive the effective Hamiltonians. In Sec. III, we investigate the generation of chiral cat states and analyze the nonclassical properties of these generated states by examining their Wigner functions. We also numerically calculate the exact state generation and verify the validity of the approximation with the fidelity. In Sec. IV, we present numerical simulations of the state generation in the open-system case and investigate the influence of the system dissipations on the state generation. Finally, we discuss the experimental implementation of this scheme and conclude this work in Sec. V.
@@ II   Physical model and Hamiltonians
We consider a cavity-QED system consisting of a two-level atom with an energy separation $\omega_{a}$ coupled to a spinning resonator supporting both the CW and CCW traveling modes via the optical evanescent fields (Fig. 1). The rotation of the resonator in the CCW direction at an angular velocity $\Omega$ induces a Sagnac-Fizeau frequency shift for both the CW and CCW modes. Then the effective frequencies of the CW and CCW modes become $\omega_{c}+\Delta_{\mathrm{sag}}$ and $\omega_{c}-\Delta_{\mathrm{sag}}$, respectively. Here, $\omega_{c}$ is the resonance frequency of the stationary resonator and $\Delta_{\mathrm{sag}}$ is the Sagnac-Fizeau frequency shift defined as [71]
@@
In Eq. (1), $\lambda$ and $c$ represent the wavelength of the field mode and the speed of light in vacuum, respectively, $n_{r}$ is the refractive index, and $R$ is the radius of the resonator. The term $dn_{r}/d\lambda$ is the dispersion term related to the relativistic origin of the Sagnac effect, which is relatively small and can be ignored in typical materials [72, 71].
@@
where $\sigma_{z}\equiv|e\rangle\langle e|-|g\rangle\langle g|$ is $z$-direction Pauli operator, with $|g\rangle$ and $|e\rangle$ being the ground and excited states of the two-level atom, respectively. The operators $\sigma_{-}\equiv|g\rangle\langle e|$ and $\sigma_{+}\equiv|e\rangle\langle g|$ are, respectively, the lowering and raising operators of the atom. The operators $a_{\epsilon}^{\dagger}$ and $a_{\epsilon}$ ($\epsilon=\mathrm{CW},\mathrm{CCW}$) are the creation and annihilation operators of the $\epsilon$ mode. The variables $\Delta_{\mathrm{CW}}\equiv\Delta+\Delta_{\mathrm{sag}}$ and $\Delta_{\mathrm{CCW}}\equiv\Delta-\Delta_{\mathrm{sag}}$ represent, respectively, the effective frequency detunings of the CW and CCW modes with respect to the resonance frequency of the atom, with $\Delta\equiv\omega_{c}-\omega_{a}$. The parameter $J$ denotes the coupling strength between the two traveling modes and the atom. We note that the bichromatic Jaynes-Cummings described by the Hamiltonian (2) can be realized in various physical systems such as coupled whispering-gallery microresonators-atom systems [80, 81, 55] and circuit QED systems [82, 83]. For the purpose of chiral state generation, we consider the coupled atom-spinning-resonator system in this work.
In this work, we consider the dispersive regime of the cavity-field-atom couplings, namely, the detunings between the two-level atom and the cavity field are much larger than the coupling strengths, i.e., $\Delta_{\epsilon=\mathrm{CW},\mathrm{CCW}}\gg J\sqrt{n_{\epsilon}+1}$ [84], where $n_{\epsilon}$ are the maximally involved photon numbers in the two cavity modes. In this case, we can perform the Frohlich-Nakajima transformation [85, 86] on the Hamiltonian (2) to derive the effective Hamiltonian by eliminating the coherent transitions between the two atomic states,
@@
$\zeta_{\epsilon=\mathrm{CW},\mathrm{CCW}}/(4\Delta_{\mathrm{sag}})\ll 1,$ (4)
@@
The physical mechanism of the chiral-cat-state generation can be understood more clearly by expressing the approximate Hamiltonian of the system in the form of conditional rotation couplings
@@
Here, the opposite signs of the coefficients for $|e\rangle\langle e|$ and $|g\rangle\langle g|$ can be used to generate cat states for the traveling modes and the different values of $\zeta_{\mathrm{CW}}$ and $\zeta_{\mathrm{CCW}}$ lead to different phase-space positions for the two traveling modes, since the effective rotating velocities of the two modes are different.
III   Generation of the chiral cat states
@@
with the normalization constants
@@
Typically, the states $|\psi_{\pm}(t)\rangle$ of the CW and CCW modes are entangled states. However, when the two coherent state amplitudes for one of the two modes are overlapped, i.e., $\alpha^{(+)}_{\mathrm{CW}/\mathrm{CCW}}=\alpha^{(-)}_{\mathrm{CW}/\mathrm{CCW}}$, then the two modes will be disentangled. For example, we set the effective detunings $\Delta_{\mathrm{CW}}=2\Delta_{\mathrm{CCW}}$; then the CW mode is in a maximally distinguishable coherent state (cat state) while the CCW mode can remain in a coherent state at time $t_{s}=\pi/(2\zeta_{\mathrm{CW}})$. We denote this phenomenon as the generation of chiral cat states, i.e., creating distinct cat states in the two different rotating modes of the cavity. Note that here the coherent state can be understood as a special case of the cat state, with a zero distance between the two superposition components in phase space, which can mostly exhibit the contrast with the cat state. The two modes of the resonator can also be prepared in two distinct cat states by choosing other measurement times for the atom. When $\Delta_{\mathrm{CW}}\neq 2\Delta_{\mathrm{CCW}}$, it becomes impossible to simultaneously achieve a coherent state and a maximally distinguishable cat state. In this case, one mode is in a coherent state, while the other mode is in a less distinguishable cat state. Nonetheless, a chiral cat state can still be generated under these conditions.
Figure 2: Wigner functions $\mathrm{W}_{\mathrm{CW,CCW}}^{(\pm)}(\chi)$ of the generated states $\rho^{(\pm)}_{\mathrm{CW,CCW}}$ at time $t_{s}=\pi/(2\zeta_{\mathrm{CW}})$. Panels (a) and (b) [(c) and (d)] show the Wigner functions $\mathrm{W}_{\mathrm{CW}}^{(+)}(\chi)$ and $\mathrm{W}_{\mathrm{CW}}^{(-)}(\chi)$ [$\mathrm{W}_{\mathrm{CCW}}^{(+)}(\chi)$ and $\mathrm{W}_{\mathrm{CCW}}^{(-)}(\chi)$] for the CW (CCW) mode, respectively. The coherent amplitude of the initial coherent state $|\alpha\rangle$ is $\alpha=1.8$. Other parameters used are $\Delta/J=33$ and $\Delta_{\mathrm{sag}}/J=11$.
To demonstrate the quantum coherence effect in the generated chiral cat states, we calculate the Wigner function [87] for both the CW and CCW modes. When the CW and CCW modes are in the states described by the density matrices $\rho_{\mathrm{CW}}$ and $\rho_{\mathrm{CCW}}$, respectively, the Wigner functions are defined by
@@
In Figs. 2(a)-2(d), we show the Wigner functions $\mathrm{W}_{\mathrm{CW,CCW}}^{(\pm)}(\chi)$ for both the CW and CCW modes at time $t_{s}=\pi/(2\zeta_{\mathrm{CW}})=69.11J^{-1}$. Concretely, Figs. 2(a) and 2(b) display the Wigner functions $\mathrm{W}_{\mathrm{CW}}^{(+)}(\chi)$ and $\mathrm{W}_{\mathrm{CW}}^{(-)}(\chi)$ for the CW mode, respectively. Here we can see distinct interference fringes caused by quantum superposition of two coherent states. The interference fringes indicate the presence of coherent superposition states, commonly referred to as cat states. In contrast, the Wigner functions for the CCW mode exhibit Gaussian peaks in the phase space, as shown in Figs. 2(c) and 2(d), which are signature for coherent states in the CCW mode. We note that the appearance of the coherent state for the CCW mode is caused by the overlap between the two coherent amplitudes $\alpha_{\mathrm{CCW}}^{(-)}(t_{s})$ and $\alpha_{\mathrm{CCW}}^{(+)}(t_{s})$. Comparing Figs. 2(a) [2(b)] with 2(c) [2(d)], we observe that the CW mode is in a cat state, while the CCW mode is in a coherent state, demonstrating the generation of chiral cat states.
@@
Figure 3: Analytical measurement probabilities $\mathcal{P}_{\pm}(t)$ (a) from Eq. (12) and the exact measurement probabilities $P_{\pm}(t)$ (b) from Eq. (19) as functions of the scaled evolution time $Jt$. For the exact solution in panel (b), both the CW and CCW modes have a truncation dimension of 13. Other parameters used in panels (a) and (b) are the same as those in Fig. 2.
@@
In Fig. 3(a), we display the analytical measurement probabilities $\mathcal{P}_{\pm}(t)$, given by Eq. (12), vs the scaled evolution time $Jt$. We observe that the sum of $\mathcal{P}_{+}(t)$ and $\mathcal{P}_{-}(t)$ is normalized. Furthermore, for longer subsequent measurement time $t_{m}$, the measurement probabilities $\mathcal{P}_{\pm}(t)\approx 1/2$, corresponding to $\sum_{l=1,-1}\exp[2li\vartheta(t_{m})+\left|\alpha\right|^{2}[\sum_{\epsilon=\mathrm{CW},\mathrm{CCW}}\mathrm{exp}(2li\zeta_{\epsilon}t_{m})-2]]\rightarrow 0$ in Eq. (12). The exact measurement probabilities $P_{\pm}(t)$ given by Eq. (19) are shown in Fig. 3(b). The exact measurement probabilities $P_{\pm}(t)$ exhibit oscillating behavior around 1/2, which is caused by the last term in Eq. (3). By comparing the analytical results with the numerical results, we find that the analytical probability evolution is smooth, while the numerical results exhibit oscillation. Nevertheless, the patterns of the approximate and exact results agree well, which confirms the validity of the approximate made during the derivation of the approximate Hamiltonian.
To quantitatively verify the validity of the approximate Hamiltonian $H_{\mathrm{app}}$, we evaluate the fidelity $F(t)=|\langle\Psi(t)|\psi_{\mathrm{app}}(t)\rangle|^{2}$ between the exact and approximate states. Here, $|\psi_{\mathrm{app}}(t)\rangle$ represents the approximate state governed by the approximate Hamiltonian $H_{\mathrm{app}}$ in Eq. (5) and $|\Psi(t)\rangle$ represents the exact state governed by the full Hamiltonian $H$ in Eq. (2). Based on Eqs. (8) and (16), the fidelity can be calculated as
@@
Similarly, we can also evaluate the performance of the chiral-cat-states generation by calculating the fidelities between the generate states $|\Psi_{\pm}(t)\rangle$ in Eq. (18) and the target states $|\psi_{\pm}(t)\rangle$ in Eq. (10). Using Eqs. (18) and (10), the fidelities $F_{\pm}(t)=|\langle\Psi_{\pm}(t)|\psi_{\pm}(t)\rangle|^{2}$ can be obtained as
@@
Figure 4: Fidelities (a) $F(t)$, (b) $F_{+}(t)$, and (c) $F_{-}(t)$ as functions of the scaled evolution time $Jt$. The parameters used are the same as those in Fig. 2.
@@
III.3 Wigner functions of the exact states
Figure 5: (a)--(d) Wigner functions $\mathcal{W}^{(\pm)}_{\mathrm{CW,CCW}}(\chi)$ for the exact states $|\Psi_{\pm}(t_{s})\rangle$ in Eq. (18) at time $t_{s}=\pi/(2\zeta_{\mathrm{CW}})$. Other parameters used are the same as those in Fig. 2.